7x=36+x+6y solve for y, step by step response
Answer by Alan3354(69443)

7x=36+x+6y
solve for y
----------
7x = 36 + x + 6y
Subtract x
6x = 36 + 6y
----------
Rewrite
6y + 36 = 6x
Subtract 36
6y = 6x - 36
Divide by 6
y = x - 6